A 4.5 kg dog stands on an 18 kg flatboat at distance D=6.1 m from the shore. It walks 2.4 m along the boat toward shore and then stops. Assuming no friction between the boat and the water, find how far the dog is then from the shore.


1
Expert's answer
2021-12-29T12:35:09-0500

Apply momentum conservation:


"0=m_bv_b-m_dv_d,\\\space\\
0=m_b\dfrac{d_b}{t}-m_d\dfrac{d_d}{t},\\\space\\
0=m_bd_b-m_dd_d."

On the other hand, these displacements added to each other give 2.4 m:


"d_b+d_d=2.4 \text{ m.}"

Solve the system:


"d_d=1.92\text{ m},\\
d_d=0.48\text{ m}."

So, dog went 1.92 m forward, and the distance to the shore became 6.1-1.92=4.18 m.